OpenStudy (anonymous):

i need help ! evaluate the integral dx/sin^2 x + cos^2 x =?

myininaya (myininaya):

what is that on the bottom?

myininaya (myininaya):

(sin^2(x)+cos^2(x))?

myininaya (myininaya):

or just sin^2(x)?

myininaya (myininaya):

because if it is (sin^2(x)+cos^2(x))? then this is really easy because sin^2(x)+cos^2(x)=1

OpenStudy (slaaibak):

I smell an identity here!

OpenStudy (anonymous):

(sin^2(x)+cos^2(x))?

OpenStudy (anonymous):

eveluate the integral \[\int\limits_{}^{} dx/\sin^2x +\cos^2x\] that s the question

myininaya (myininaya):

\[\int\limits_{}^{}(\frac{1}{\sin^2(x)} +\cos^2(x)) dx\]

myininaya (myininaya):

is that right?

myininaya (myininaya):

your problem is lacking some parenthesis

OpenStudy (anonymous):

\[\int\limits_{}^{} 1/(\sin^2x + \cos^2x)dx\] that one true

myininaya (myininaya):

i was assuming at first you meant: \[\int\limits_{}^{}\frac{dx}{\sin^2(x)+\cos^2(x)} \]

OpenStudy (anonymous):

oh yes lol

myininaya (myininaya):

did you not see what i said above if this what you meant?

myininaya (myininaya):

\[\sin^2(x)+\cos^2(x)=1\]

myininaya (myininaya):

\[\int\limits_{}^{}1 dx=x+C\]

OpenStudy (anonymous):

oh sorry sorry sin^2 (x) +sinx at the bottom

OpenStudy (anonymous):

i reaaly confused

myininaya (myininaya):

i accidently pressed the back button :(

OpenStudy (anonymous):

did u solve ?

myininaya (myininaya):

i will do it on paper can't hit the back button there lol yeah i was almost done

myininaya (myininaya):

so just to be sure i'm doing \[\int\limits_{}^{}\frac{dx}{\sin^2(x)+\sin(x)}\]

OpenStudy (anonymous):

lol it s ok. i m waiting...yes be sure. lol i am new member to this site and i don t know many things yet lol

myininaya (myininaya):

ok done most of it i left one thing for you

myininaya (myininaya):